OpenStudy (dawnr):

from all the points of the circle x^2+y^2=4 point(x0,y0) is the farthest from the line: x-2y-1=0 so x0+y0=?

OpenStudy (anonymous):



OpenStudy (anonymous):

from the line, we have y = 1/2x -1 so that the perpendicular line to that line has the form y=-2x +C

OpenStudy (anonymous):



OpenStudy (dawnr):

okay..

OpenStudy (anonymous):

the line will cut the circle at 2 points and the farthest point on the circle will be the perpendicular line with the given line at the middle point, right?

OpenStudy (dawnr):

right!

OpenStudy (anonymous):

and every line perpendicular to the arc at the middle point passes through the center, right?

OpenStudy (anonymous):

you have the center, you have the middle point, you have the form of the circle, you can find x0, y0, do whatever you want

OpenStudy (dawnr):

but how do i know which one to use for what?! :O

OpenStudy (anonymous):

the perpendicular line pass through the center is y =-2x (because y =0) (1) so that , if this line cut the circle, you will have y^2 +x^2 =4 (2) square both sides of (1) you have y^2 =4x^2, replace to (2) you have 5x^2=4, x =\(\pm\dfrac{2}{\sqrt5}\) , replace back to (1) to find y

OpenStudy (dawnr):

another question how do i know if i should use + or - value of x?

OpenStudy (dawnr):

when i sum those two..

OpenStudy (anonymous):

As I told before, when you get x0, y0, do whatever you want, times them, add them, subtract them, exponent them and whatever....... You can "eat" them too. hihihih

OpenStudy (dawnr):

XD yea but the answer isn't the same :D

OpenStudy (anonymous):

sure, because when you eat them, they disappear, hihihiih... Whatsoever, you have the way to find them out. Now is your duty. me done

OpenStudy (dawnr):

okay :p

ganeshie8 (ganeshie8):

nice :)